An electrical appliance works with a voltage not exceeding 25V when it is connected to the main power supply a step-down transformer with 50 turns in the primary coil is used.Calculate the number of turns in the secondary coil when the voltage of the main power supply is 220V.


1
Expert's answer
2022-06-14T08:41:01-0400
"\frac{V_1}{V_2}=\frac{N_1}{N_2}"

"N_2=N_1\frac{V_2}{V_1}=50*\frac{220}{25}=440\:\rm turns"